find a polynomial for the perimeter and area of a rectangle when the long side is d+7 and the short side is d.
perimeter=?
area=?

Answer by mananth(16946)

width = d
length =d+7
perimeter = 2L+2D
2(d+7)+2d
2d+14+2d
=4d+14 = perimeter
Area = L*D
=d*(d+7)
=d^2 +7d = area